OpenStudy (anonymous):

help me $6,000 dollars is invested in two different accounts earning 3% and 5% interest. At the end of one year, the two accounts earned $220 in interest. How much money was invested at 3%? $2,000 $3,000 $4,000

ganeshie8 (ganeshie8):

lets say, amount invested in first account x amount invested in second account y x + y = 6000

ganeshie8 (ganeshie8):

interest on x = 3% = .03x interest on y = 5% = .05y since, total interest earned is 220 .03x + .05y = 220 3x + 5y = 22000

ganeshie8 (ganeshie8):

so, we got two equations : x + y = 6000 3x + 2y = 22000

ganeshie8 (ganeshie8):

can you solve for x now ?

OpenStudy (anonymous):

i had got 2000

ganeshie8 (ganeshie8):

wait a sec. . let me check

ganeshie8 (ganeshie8):

i am getting 4000 in first account(account with 3% interest rate) & 2000 in second account.

OpenStudy (anonymous):

so was i right

ganeshie8 (ganeshie8):

nope $4000 is the answer

OpenStudy (anonymous):

ook thx

ganeshie8 (ganeshie8):

yw